OpenStudy (anonymous):

13. A skier, starting from rest, slides down a frictionless surface with a 30 degree incline from an initial height of 180 m. How fast is she going when she reaches the bottom of the slope? Use energy methods to solve.

OpenStudy (anonymous):

from the energy conservation: total mechanical energy at the top=total mechanical energy at the down 0(K.E)+mgh=1/2mv^2+0 v=sqrt(2gh)=sqrt(2*9.8*180)=60m/sec